APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday October 15, 2026 (SJD 61329) ---- Observing Team ---- Kaike Pan Jose Sanchez-Gallego (SOS Page) Shane Tomas (Lui and apogeeql reduction page) John Donor (apogeeql) ---- Observing Plan ---- Thu 15 Oct 2026 61329 0.28 01:23 01:37 12:00 20:12 06:37 --- dark (10.4 h) plus 0.2 h for 12 deg design ---- Observing Summary ---- High wind speed, between 35 to 65 mph, in the entire night. We never open. Notes: 1. There was a power glitch during a test slew from instrument-change to stow in afternoon check, which caused axes went out-of-closed-loop. After the power recover (in one second), "axis init" cleared all axes normally, but altitude axis went out-of-closed-loop after less than 1 deg motion. Three attempts gave the same result until I engaged/released an e-stop, the slew completed normally. ---- Observing Log ---- Afternoon: ========== After server maintenance, several webpage on sdss5-hub give permission error. Jose and Shane fixed. John Donor fixed the apogeeql issues(kept dropping from STUI and didn't bundle data). BOSS collimation checked, no adjustment made. 3:20:11 Hartmann collimate output sp1: offset: r= -659 (In Focus); b= -2.2 (In Focus) sp1: pred. move: spAverageMove= 24 sp1: pred. resid: r= -683, b= -2.1, txt= OK, spTemp = 7.0 Although APOGEE takes data normally and saves the raw data to apogee-ics computer, but apogeeql doesn't bundle, archive the data correct. Because John was not available until later, I went head to take the first set of apogee evening cal. Wind speed is up to 50 mph, after fixed apogeeql issues, I took 2nd set of apogee evening cal. BOSS cal, apogee evening cal, fvc dark, GFA biases, dark, flat were taken, checked. Cherno restarted. High wind prevents us to open at the sunset, delayed open. High wind speed, we never open. ---- Mirror Numbers ---- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: -5.4290 -5.4260 Axial B: -5.0400 -5.0390 Axial C: -5.2820 -5.2780 Trans D: -0.8760 -0.8710 Lateral E: 0.3240 0.3245 Lateral F: 0.5335 0.5335 GALILS Commanded: 11450 -1050 5900 26550 -14400 -13850 Actual: 11450 -1050 5900 26550 -14400 -13850 SETMIR VALUES PrimDesOrient: 0.0 -43.05 -8.77 -751.82 -556.2 PrimOrient: -1.06 -43.05 -8.77 -751.82 -556.2 SECONDARY: ---------- Focus: 0.0000 Air temp: 0.00 Altitude: 30.0 MIGS TONIGHT NOMINAL Axial A: -7.9560 -7.9570 Axial B: -6.1140 -6.1140 Axial C: -6.0760 -6.0790 Trans D: 0.1980 0.2050 GALILS Commanded: 362900 384850 343200 -1500 1800 Actual: 321895 350059 299495 -1500 1800 SETMIR VALUES SecDesOrient: 257.0 14.2 5.9 -72.0 -12.76 SecOrient: 256.98 14.24 5.9 -71.78 -12.65 ---- Telescope Status ---- N/A Started ventilation N/A encl25m : open N/A encl25m : closed No interlocks bypassed Status at: 12:30Z Telescope stowed at: 121.0, 30.0, 0.0 mount Instrument mounted: FPS Counterweights at: LN2 autofill systems: Connected and turned on 180L LN2 dewar scales: SP1 137 lbs, 24.1 psi ---- Software Used ---- -- STUI software: STUI : 1.7.21 alerts : 2.3.0 boss : 6.0.5a1 -boss.specMec : 3.0.4 apo : 4.1.0 fliswarm : 0.7.0 jaeger : 1.11.2a1 kaiju : 1.4.1 coordio : 1.18.1 fps_calibrati : 2025.12.04+3e7e2d8 cherno : 1.2.2 astrometrynet : 0.97 hal : 1.4.1 hartmann : 2.0.2 hub : 5.0.1a0 mcp : 2.3.1 -mcp.plc : 35 -mcp.azFiduci : v1_154 -mcp.altFiduc : v1_154 -mcp.rotFiduc : v1_154 tcc : 1.4.0 apogee : not availble apogeecal : not availble apogeeql : 2.0.0a0 STUIAdditions : FAILED Boson : 0.7.0 sdss-obstools : 3.9.3 TPM : 1.2-31-gd7eb MCP : 2.3.1 PLC : V35 Fiducials : v1_154 idlspec2d : 6.3.0dev Kronos : 2.1.0a roboscheduler : 2.1.0 ---- Time Tracking ---- *=*=*= Observatory: APO *=*=*= MJD: 61329 Fill night time in Hours below *=*=*= Total bright time : 0.2 *=*=*= Bright time, queued observations : 0 *=*=*= Bright time, commissioning/engineering: 0 *=*=*= Bright time lost to weather : 0.2 *=*=*= Bright time lost to technical issues : 0 *=*=*= Total dark time : 10.4 *=*=*= Dark time, queued observations : 0 *=*=*= Dark time, commissioning/engineering : 0 *=*=*= Dark time lost to weather : 10.4 *=*=*= Dark time lost to technical issues : 0